40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 6 14:26:52 2021, Koji, Update, VAC, Pumping down the RGA section 
    Reply  Tue Dec 7 13:12:50 2021, Koji, Update, VAC, Pumping down the RGA section 
       Reply  Wed Dec 8 10:14:43 2021, Jordan, Update, VAC, Pumping down the RGA section 40m_RGAVolume_12_8_21.PNG40m_RGAVolume_Torr_12_8_21.PNG
          Reply  Fri Dec 10 19:22:01 2021, Koji, Update, VAC, Pumping down the RGA section 
Message ID: 16501     Entry time: Fri Dec 10 19:22:01 2021     In reply to: 16494
Author: Koji 
Type: Update 
Category: VAC 
Subject: Pumping down the RGA section 

The scan result was ~x10 higher than the previously reported scan on 2020/9/15 (https://nodus.ligo.caltech.edu:8081/40m/15570), which was sort of high from the reference taken on 2018/7/18.

This just could mean that the vacuum level at the RGA was x10 high.
We'll just go ahead with the vacuum repair and come back to the RGA once we return to "vacuum normal".

Meanwhile, I asked Jordan to turn off the RGA to make it cool down. I shut off RGA section and turned TP2 off.

ELOG V3.1.3-